# On independence complexes of graph products

### v1

@@ Abstract
We study the independence complex for graph products where at least one of the factors is a path as well as the complex es for their induced subgraphs. We calculate the homotopy type of the independence complex of the graphs $P_{n}\times P_{m}$, $P_{n}\boxtimes P_{2}$, $P_{n}\boxtimes P_{3}$ and $P_{n}\boxtimes P_{4}$. We also focus in the independence complexes of the induced subgraphs of $P_{n}\times P_{3}$, $P_{n}\boxtimes P_{2}$, $P_{n}\boxtimes P_{3}$, $P_{n}\boxtimes P_{4}$ and some lexicographic products.
@@ 1 Introduction
The independence complex is one of the most studied graph complexes in topological combinatorics, it is define as the set of independent sets of a graph. This complex has many applications in other branches in mathematics, for example it has been used to study the extreme Khovanov (co)homology of link diagrams (see [15, 26, 27]). Another way to think the independence complex is to define it as the clique complex of the complement graph, i.e. the simplices are the complete graphs of the complement graph; in a finite metric space the Vietoris-Rips complexes are clique complexes (see for example [2, 3, 6]) and this complexes are used in topological data analysis and geometric group theory (see for example [25]).
In general is not easy to determine the homotopy type of the independence complex. Even for highly symmetric graphs the topology of this complex can be really hard to understand, for example the chessboard complex– which can be seen as the independence complex of the cartesian product of complete graphs– is not well understood and has torsion for many cases (see [14, 18, 29, 30, 31]). For this reason much of the work is concentrated in calculating the homotopy type for specific families of graphs (see [1, 7, 8, 16, 17, 20]) and for graph products of particular graphs (see for example 4, 9, 21, 22, 23). Here we will focus in study the independence complex for graph products where at least one of the factors is a path as well as the complexes for their induced subgraphs. The products $P_{n}\boxtimes P_{m}$ will be our main objective, but we will do some calculations for categorical and lexicographic products.
@@ 2 Preliminaries
A simplicial complex $K$ is a family of subsets of an finite set $V(K)$, the vertices of the complex, such that if $\tau\subseteq\sigma$ and $\sigma\in K$, then $\tau\in K$. We will not distinguish between a simplicial complex and its geometric realization.
We are only interested in simple graphs. For a graph $G$, we denote its vertex set as $V(G)$ and its edge set as $E(G)$, the order of the graph is the non-negative integer number $|V(G)|$. For a vertex $v$, $N_{G}(v)=\{u\in V(G):\;uv\in E(G)\}$ is its open neighborhood and $N_{G}[v]=N_{G}(G)\cup\{v\}$ its closed neighborhood. The degree of a vertex will be denoted $d_{G}(v)=|N_{G}(v)|$.
@@
For two vertex disjoint graphs $G$ and $H$ their union is the graph $G+H$ with vertex set $V(G)\cup V(H)$ and edge set $E(G)\cup E(H)$. Given two graphs $G$ and $H$, we will consider four different graphs over the vertex set $V(G)\times V(H)$:
@@
For all the graph definitions not stated here we follow [10].
A vertex set $\sigma\subseteq V(G)$ is independent if any two vertices $u,v$ do not form an edge in $G$. The independence complex of $G$ is the simplicial complex
$I(G)=\{\sigma\subseteq V(G):\;\sigma\mbox{ is independent}\}$
@@
    [13] If $N_{G}(u)\subseteq N_{G}(v)$, then $I(G)\simeq I(G-v)$.
@@
    [26] If $N_{G}[u]\subseteq N_{G}[v]$, then $I(G)\simeq I(G-v)\vee\Sigma I(G-N_{G}[v])$.
@@
[9] Let $G$ be a graph with $v$ a vertex of degree $2$ such that $N_{G}(v)=\{v_{1},v_{2}\}$ is an independent set. We take the graph $H$ obtain from $G$ by first erasing $N_{G}(v_{1})\cap N_{G}(v_{2})$ from $G$, and next adding the edges $u_{1}u_{2}$ for all $u_{1}\in N_{G}(v_{1})-N_{G}(v_{2})$ and $u_{2}\in N_{G}(v_{2})-N_{G}(v_{1})$. Then
$I(G-N_{G}(v_{1}))\cup I(G-N_{G}(v_{2}))=I(H)$
$I(G-N_{G}(v))\cap\left(I(G-N_{G}(v_{1}))\cup I(G-N_{G}(v_{2}))\right)=I(H-v_{1}-v_{2})$
@@
[5] Let $G$ be a graph and let $v$ be a non-isolated vertex of $G$ which is contained in no triangle. Then
$I(G)\simeq\Sigma(I(G-N_{G}(v))\cap SC(v))$
@@
[11] Let $G$ be a graph and $uv$ an edge. Take $H$ the graph obtained from $G$ by replacing the edge $uv$ by a path of length $4$. Then $I(H)\simeq\Sigma I(G)$.
@@
[20]
$I(P_{n})\simeq\left\{\begin{array}[]{cc}\mathbb{S}^{k-1}&\mbox{if }n=3k\\ \ast&\mbox{if }n=3k+1\\ \mathbb{S}^{k}&\mbox{if }n=3k+2\end{array}\right.,\;\;I(C_{n})\simeq\left\{\begin{array}[]{cc}\displaystyle\bigvee_{2}\mathbb{S}^{k-1}&\mbox{if }n=3k\\ \mathbb{S}^{k-1}&\mbox{if }n=3k+1\mbox{ or }n=3k-1\end{array}\right.$
Before ending this section we will define two graph families. We will say a graph $G$ has property $\mathcal{P}$ if $I(G)$ is contractible, has the homotopy type of a wedge of spheres or has the homotopy type of a disjoint union of wedges of spheres. We will say a graph $G$ has property $\mathcal{SP}$ if $I(G)$ is contractible or has the homotopy type of a wedge of spheres. From this, we define the following two families:
$\mathcal{W}=\{G:\mbox{ any induced subgraph of }G\mbox{ has property }\mathcal{P}\}$
$\mathcal{SW}=\{G:\mbox{ any induced subgraph of }G\mbox{ has property }\mathcal{SP}\}$
For example, by Theorem 6, all the paths and cycles are in $\mathcal{SW}$. By the definition of the independence complex is clear that complete graphs are also in $\mathcal{SW}$. It is known that the independence complex of a forest is eihter contractible or has the homotopy type of a sphere [12, see Corollary 6.1], thus all forest are also in $\mathcal{SW}$. All graphs without induced cycles of length divisible by $3$ are also in $\mathcal{SW}$ [19].
3 Independence complexes of categorical
In this section we give a decomposition formula for the homotopy type of the independence complex of $G\times P_{n}$ in terms of joins of the complex $I(G\times P_{2})$. With this we calculate the homotopy type of $I(P_{n}\times P_{m})$. We begin giving a Lemma that was made as a commentary in [4] after the Proposition 3.1 of that article.
@@ 3 Independence complexes of categorical
$I(G\times H)\simeq I(G\times(H-u))$
@@
$N_{G\times H}((w,v))=N_{G}(w)\times N_{H}(v)\subseteq N_{G\times H}((w,u))=N_{G}(w)\times N_{H}(u)$
By Lemma 1 we can erase all the vertices $(w,u)$ without changing the homotopy type. ∎
@@
$I(G\times C_{4})\simeq I(G\times P_{2})$
@@
$I(G\times P_{n})\simeq\left\{\begin{array}[]{cc}I(G\times P_{2})^{*r}&\mbox{ if }n=3r\\ &\mbox{ if }n=3r+1\\ I(G\times P_{2})^{*r+1}&\mbox{ if }n=3r+2\end{array}\right.$
@@
For $n=1,2$, the result is clear. For $n\geq 3$, By Lemma 7 we have that
@@
The results follows by induction. ∎
@@
$I(G\times P_{n})\simeq\left\{\begin{array}[]{cc}I(G)^{*2r}&\mbox{ if }n=3r\\ &\mbox{ if }n=3r+1\\ I(G)^{*2r+2}&\mbox{ if }n=3r+2\end{array}\right.$
@@
Because $G$ is bipartite, we have that $G\times P_{2}\cong 2G$. ∎
As immediate consequence we get the following corollary.
@@
Let $G$ be a bipartite graph in $\mathcal{W}$. Then $G\times P_{2}$, then $I(W)$ is in $\mathcal{SW}$.
Proposition 12.
$I(P_{n}\times P_{m})\simeq\left\{\begin{array}[]{cc}\displaystyle\mathbb{S}^{2kr-1}&\mbox{if }n=3k\mbox{ and }m=3r\\ \displaystyle\ast&\mbox{ if }n=3k+1\mbox{ or }m=3r+1\\ \displaystyle\mathbb{S}^{2(k+1)r-1}&\mbox{if }n=3k+2\mbox{ and }m=3r\\ \displaystyle\mathbb{S}^{2(r+1)k-1}&\mbox{if }n=3k\mbox{ and }m=3r+2\\ \displaystyle\mathbb{S}^{2(k+1)(r+1)-1}&\mbox{if }n=3k+2\mbox{ and }m=3r+2\end{array}\right.$
Proof.
Thus result follows from Theorem 6 and Corollary 10. ∎
In [4, Proposition 3.4.] was proved that
$I(C_{m}\times P_{2})\simeq\left\{\begin{array}[]{cc}\displaystyle\bigvee_{4}\mathbb{S}^{4k-1}&\mbox{if }m=6k\\ \mathbb{S}^{4k}&\mbox{if }m=6k+1\\ \mathbb{S}^{4k+1}&\mbox{if }m=6k+r\;\mbox{with }r\in\{2,4\}\\ \mathbb{S}^{4k+1}\vee\mathbb{S}^{4k+1}&\mbox{if }m=6k+3\\ \mathbb{S}^{4k+2}&\mbox{if }m=6k+5\end{array}\right.$
this together with Theorem 9 give us the following corollary.
Corollary 13.
$I(C_{m}\times P_{n})\simeq\left\{\begin{array}[]{cc}\ast&\mbox{ if }n=3r+1\\ \displaystyle\bigvee_{4^{r}}\mathbb{S}^{4kr-1}&\mbox{if }m=6k\mbox{ and }n=3r\\ \mathbb{S}^{4kr+r-1}&\mbox{if }m=6k+1\mbox{ and }n=3r\\ \mathbb{S}^{4kr+2r-1}&\mbox{if }m=6k+l\;\mbox{with }l\in\{2,4\},\mbox{ and }n=3r\\ \displaystyle\bigvee_{2^{r}}\mathbb{S}^{4kr+2r-1}&\mbox{if }m=6k+3\mbox{ and }n=3r\\ \mathbb{S}^{4kr+3r-1}&\mbox{if }m=6k+5\mbox{ and }n=3r\\ \displaystyle\bigvee_{4^{r+1}}\mathbb{S}^{4k(r+1)-1}&\mbox{if }m=6k\mbox{ and }n=3r+2\\ \mathbb{S}^{4k(r+1)+r}&\mbox{if }m=6k+1\mbox{ and }n=3r+2\\ \mathbb{S}^{4k(r+1)+2r+1}&\mbox{if }m=6k+l\;\mbox{with }l\in\{2,4\},\mbox{ and}n=3r+2\\ \displaystyle\bigvee_{4^{r+1}}\mathbb{S}^{4k(r+1)+2r+1}&\mbox{if }m=6k+3\mbox{and }n=3r+2\\ \mathbb{S}^{4k(r+1)+3r+2}&\mbox{if }m=6k+5\mbox{ and }n=3r+2\end{array}\right.$
@@
$\mathrm{conn}(I(G\times P_{3r}))\geq r+1\;\;\mbox{and}\;\;\mathrm{conn}(I(G\times P_{3r+2}))\geq r+2.$
@@
For $n=1$ the result is clear. If we prove the result is for $n=2$, then, by Theorem 9, the result is true for any $n\geq 3$. Assume $n=2$ and let $T$ be a spanning tree of $G^{c}$. We define $T_{i}$ as the copy of $T$ in $\Delta^{V(G)\times\{i\}}$ and we take $x$ a vertex in $V(G)$. We define
@@
Then,
$\pi_{1}\left(I(G\times P_{2}),(x,1)\right)\cong H_{\mathcal{T}}$
(see [28, Theorem 7.34]). We will see that $H_{\mathcal{T}}$ is trivial. Take $e$ an edge of $(G\times P_{2})^{c}$:
@@
  • If $e=\{(w,1),(w,2)\}$, we take $y_{0}y_{1}\cdots y_{r}$ the only path form $w$ to $x$ in $T_{1}$. For all $i$, we have that
@@
the rest of possibilities are analogous. Therefore $H_{\mathcal{T}}$ is trivial.
The result follows form what we have done and Theorem 9. ∎
The lower bound given in last Theorem is tight, to see this take $C_{5}\times P_{2}\cong C_{10}$, by Theorem 6 $I(C_{5}\times P_{2})\simeq\mathbb{S}^{2}$.
@@
For $m=1$, we have that $P_{n}\times P_{1}\cong K_{n}^{c}$ and the result is clear. For $m=2$, we have that $P_{n}\times P_{2}\cong 2P_{n}$ and any induced subgraph is the disjoint union of paths.
For $P_{n}\times P_{3}$ we know it is true for $n\leq 2$. Assume the result is true for $n$ and take $H$ an induced subgraph of $P_{n+1}\times P_{3}$. If neither of the vertices $(1,1),(1,2),(1,3)$ are in $V(H)$, then $H$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{3}$ and there is nothing to prove.
@@
Regardless of the case, we got that $I(H)$ is either contractible or has the homotopy type of the double suspension of a complex $I(H^{\prime})$ where $H^{\prime}$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{3}$. If $(1,3)$ is in $H$, we proceed as before. If $(1,2)$ is a vertex of $H$ and $(1,1),(1,3)$ are not in $V(H)$, we proceed as the last case when we assumed $(1,2)$ was a vertex of $H-N_{H}[(2,2)]$.
Lastly, assume $m=4$. We know that the result is true for $n\leq 3$. Assume the result is true for $n$ and take $H$ and induced subgraph of $P_{n+1}\times P_{3}$. If neither of the vertices $(1,1),(1,2),(1,3),(1,4)$ are in $V(H)$, then $H$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{3}$ and there is nothing to prove. If the vertex $(1,1)$ is in $H$, it has at most degree $1$, then $I(H)$ is contractible or $I(H)\simeq\Sigma I(N_{H}[(2,2)])$. If $(1,4)$ is in $H$, then $d_{H-N_{H}[(2,2)]}((1,4))=d_{H}((1,4))\leq 1$ and $I(H-N_{H}[(2,2)])$ is contractible or
@@
with $H-N_{H}[(2,2)]-N_{H}[(2,3)]$ an induced subgraph of a graph isomorphic to $P_{n}\times P_{4}$. Assume $(1,4)$ is not in $H$. If $(2,4)$ is a vertex of $H$, then $I(H-N_{H}[(2,2)])$ is contractible. Assume $(2,4)$ is not in $H$. If $(1,2)$ is not in $H$, then $H-N_{H}[(2,2)]$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{4}$. Assume $(1,2)$ is a vertex of $H$, then $d_{H-N_{H}[(2,2)]}((1,2))=d_{H}((1,2))\leq 2$. If $d_{H}((1,2))=0$, then $I(H-N_{H}[(2,2)])$ is contractible. If $d_{H}((1,2))=1$, then
@@
Assume $(3,2)$ is a vertex of $H$, then
@@
In either case $I(H)$ is contractible or has the homotopy type of the double suspension of a complex $I(H^{\prime})$ where $H^{\prime}$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{4}$. If $(1,4)$ is in $H$, we proceed as before. If $(1,2)$ or $(1,3)$ are vertices of $H$ and $(1,1),(1,4)$ are not in $V(H)$, we proceed as the last case when we assumed $(1,2)$ was a vertex of $H-N_{H}[(2,2)]$. ∎
Question 16.
@@
Before ending this section we want to remark that while the homotopy type of the independence complex of the categorical product of paths was easy to determine, the homotopy type of the independence complexes of $P_{n}\oblong P_{m}$ is only know for $1\leq m\leq 6$ (see [21, 22]).

### v2

@@ Abstract
We study the independence complexes of graph products where at least one factor is a path. We also analyze the complexes of their induced subgraphs. We determine the homotopy type of the independence complex of the graphs $P_{n}\times P_{m}$, $P_{n}\boxtimes P_{2}$, $P_{n}\boxtimes P_{3}$ and $P_{n}\boxtimes P_{4}$. We also focus in the independence complexes of the induced subgraphs of $P_{n}\times P_{3}$, $P_{n}\boxtimes P_{2}$, $P_{n}\boxtimes P_{3}$, $P_{n}\boxtimes P_{4}$ and some lexicographic products $G\circ H$.
@@ 1 Introduction
The independence complex is one of the most studied graph complexes in topological combinatorics, it is define as the set of independent sets of a graph. This complex has many applications in other branches in mathematics, for example it has been used to study the extreme Khovanov (co)homology of link diagrams (see [Gonz_lez_Meneses_2017, marithaniacircindcompl, MR4741186]). Another way to think the independence complex is to define it as the clique complex of the complement graph, i.e. the simplices are the complete graphs of the complement graph; in a finite metric space the Vietoris-Rips complexes are clique complexes (see for example [adamaviet, nervcompcirarc, bendersky2023connectivity]) and this complexes are used in topological data analysis and geometric group theory (see for example [toppersgeo]).
In general is not easy to determine the homotopy type of the independence complex. Even for highly symmetric graphs the topology of this complex can be really hard to understand, for example the chessboard complex– which can be seen as the independence complex of the cartesian product of complete graphs– is not well understood and has torsion for many cases (see [Jonsson_2010, Shareshian_2007]). For this reason much of the work is concentrated in determining its homotopy type for specific families of graphs (see [adamsplit, Bousquet_M_lou_2007, Braun_2017, Iriye_2012, Jonsson_2009, kozlovdire]) and graph products of particular graphs (see for example indcomplcatprod, poljoin, matsushitan4n5, matsushitan6, MR4477852). Here we will focus in study the independence complexes of graph products in which one factor is a path, along with those of their induced subgraphs. Our main focus will be the complexes of $P_{n}\boxtimes P_{m}$, but we will do some calculations for categorical products and lexicographic products.
@@ 2 Preliminaries
A simplicial complex $K$ is a family of subsets of a finite set $V(K)$, the vertices of the complex, such that if $\tau\subseteq\sigma$ and $\sigma\in K$, then $\tau\in K$. We will not distinguish between a simplicial complex and its geometric realization.
We are only interested in simple graphs. For a graph $G$, we denote its vertex set as $V(G)$ and its edge set as $E(G)$. The order of a graph $G$ is the non-negative integer number $|V(G)|$. For a vertex $v$, $N_{G}(v)=\{u\in V(G):\;uv\in E(G)\}$ is its open neighborhood and $N_{G}[v]=N_{G}(G)\cup\{v\}$ its closed neighborhood. The degree of a vertex will be denoted $d_{G}(v)=|N_{G}(v)|$.
@@
For two vertex disjoint graphs $G$ and $H$ their disjoint union is the graph $G+H$ with vertex set $V(G)\cup V(H)$ and edge set $E(G)\cup E(H)$. Given two graphs $G$ and $H$, we will consider four different graphs over the vertex set $V(G)\times V(H)$:
@@
For all the graph definitions not stated here we follow [graphsanddigraphs].
A vertex set $\sigma\subseteq V(G)$ is independent if no two vertices in $\sigma$ are adjacent. The independence complex of $G$ is the simplicial complex
$I(G)=\{\sigma\subseteq V(G):\;\sigma\mbox{ is independent}\}.$
@@
    [engstrom09] If $N_{G}(u)\subseteq N_{G}(v)$, then $I(G)\simeq I(G-v)$.
@@
    [marithaniacircindcompl] If $N_{G}[u]\subseteq N_{G}[v]$, then $I(G)\simeq I(G-v)\vee\Sigma I(G-N_{G}[v])$.
@@
[poljoin] Let $G$ be a graph with $v$ a vertex of degree $2$ such that $N_{G}(v)=\{v_{1},v_{2}\}$ is an independent set. We take $H$ the graph obtain from $G$ by deleting the vertices in $N_{G}(v_{1})\cap N_{G}(v_{2})$ from $G$ and adding the edges $u_{1}u_{2}$ for all $u_{1}\in N_{G}(v_{1})-N_{G}(v_{2})$ and $u_{2}\in N_{G}(v_{2})-N_{G}(v_{1})$. Then
$I(G-N_{G}(v_{1}))\cup I(G-N_{G}(v_{2}))=I(H),$
$I(G-N_{G}(v))\cap\left(I(G-N_{G}(v_{1}))\cup I(G-N_{G}(v_{2}))\right)=I(H-v_{1}-v_{2}).$
@@
[barmak] Let $G$ be a graph and let $v$ be a non-isolated vertex of $G$ which is contained in no triangle. Then
$I(G)\simeq\Sigma(I(G-N_{G}(v))\cap SC(v)).$
@@
[csorba] Let $G$ be a graph and $uv$ an edge. Take $H$ the graph obtained from $G$ by replacing the edge $uv$ by a path of length $4$. Then $I(H)\simeq\Sigma I(G)$.
@@
[kozlovdire]
$I(P_{n})\simeq\left\{\begin{array}[]{cc}\mathbb{S}^{k-1}&\mbox{if }n=3k\\ \ast&\mbox{if }n=3k+1\\ \mathbb{S}^{k}&\mbox{if }n=3k+2\end{array}\right.,\;\;I(C_{n})\simeq\left\{\begin{array}[]{cc}\displaystyle\bigvee_{2}\mathbb{S}^{k-1}&\mbox{if }n=3k\\ \mathbb{S}^{k-1}&\mbox{if }n=3k+1\mbox{ or }n=3k-1\end{array}\right.$.
Before we conclude this section we define two families of graphs. We say a graph $G$ has property $\mathcal{P}$ if either $I(G)$ is contractible, it has the homotopy type of a wedge of spheres or it has the homotopy type of a disjoint union of wedges of spheres. We say a graph $G$ has property $\mathcal{SP}$ if $I(G)$ is either contractible or it has the homotopy type of a wedge of spheres. From this, we define the following two families:
[…]
For example, by Theorem 6, all the paths and cycles are in $\mathcal{SW}$. By the definition of the independence complex is clear that complete graphs are also in $\mathcal{SW}$. It is known that the independence complex of a forest is either contractible or has the homotopy type of a sphere [Ehrenborg_2006, see Corollary 6.1], thus all forests are also in $\mathcal{SW}$. Lastly all graphs without induced cycles of length divisible by $3$ have independence complex homotopy equivalent to a sphere or contractible [kimternary], thus these graphs are also in $\mathcal{SW}$.
Proposition 7.
If $G_{1},\dots,G_{k}$ are in $\mathcal{W}$ or in $\mathcal{SW}$, then $G_{1}+\cdots+G_{k}$ are in $\mathcal{W}$ or in $\mathcal{SW}$ respectively.
Proof.
We take $H$ an induced subgraph of $G_{1}+\cdots+G_{k}$. Let $A=\{i_{1},\dots,i_{r}\}$ the subset of $\underline{k}$ such that $H$ and $G_{i_{j}}$ have at least a common vertex for any $i_{j}$ in $A$. Then, for each $i_{j}$ in $A$ there is $H_{i_{j}}$ an induced subgraph of $G_{i_{j}}$ such that
$H=H_{i_{1}}+\dots+H_{i_{r}}.$
Thus $I(H)=I(G_{i_{1}})*\dots*I(H_{i_{r}})$. ∎
3 Independence complexes of categorical products
In this section we give a decomposition formula for the homotopy type of the independence complex of $G\times P_{n}$ in terms of joins of the complex $I(G\times P_{2})$. With this we calculate the homotopy type of $I(P_{n}\times P_{m})$. We begin giving a Lemma that was made as a commentary in [indcomplcatprod] after the Proposition 3.1 of that article.
@@ 3 Independence complexes of categorical products
$I(G\times H)\simeq I(G\times(H-u)).$
@@
$N_{G\times H}((w,v))=N_{G}(w)\times N_{H}(v)\subseteq N_{G\times H}((w,u))=N_{G}(w)\times N_{H}(u).$
By Lemma 1 we can delete all the vertices $(w,u)$ without changing the homotopy type. ∎
@@
$I(G\times C_{4})\simeq I(G\times P_{2}).$
@@
$I(G\times P_{n})\simeq\left\{\begin{array}[]{cc}I(G\times P_{2})^{*r}&\mbox{ if }n=3r\\ &\mbox{ if }n=3r+1\\ I(G\times P_{2})^{*r+1}&\mbox{ if }n=3r+2\end{array}\right.$.
@@
The proof is by induction on $n$. For $n=1,2$ this is clear. For $n\geq 3$, By Lemma 8 we have that
@@
Thus $I(G\times P_{3})\simeq I(G\times P_{2})$. The rest follows by induction. ∎
If a graph $G$ is bipartite, then is easy to see that $G\times P_{2}\cong 2G$. From this we get the following two corollaries.
@@
$I(G\times P_{n})\simeq\left\{\begin{array}[]{cc}I(G)^{*2r}&\mbox{ if }n=3r\\ &\mbox{ if }n=3r+1\\ I(G)^{*2r+2}&\mbox{ if }n=3r+2\end{array}\right.$.
Corollary 12.
Let $G$ be a bipartite graph in $\mathcal{W}$. Then $G\times P_{2}$ is in $\mathcal{W}$.
Question 13.
If $G$ is a bipartite graph in $\mathcal{W}$, Does $G\times P_{n}$ is in $\mathcal{W}$ for all $n$?
Now we can easily determine the homotopy type of $I(P_{n}\times P_{m})$.
Proposition 14.
[…]
@@
The proposition follows from Theorem 6 and Corollary 11. ∎
In [indcomplcatprod, Proposition 3.4.] it was proved that
$I(C_{m}\times P_{2})\simeq\left\{\begin{array}[]{cc}\displaystyle\bigvee_{4}\mathbb{S}^{4k-1}&\mbox{if }m=6k\\ \mathbb{S}^{4k}&\mbox{if }m=6k+1\\ \mathbb{S}^{4k+1}&\mbox{if }m=6k+r\;\mbox{with }r\in\{2,4\}\\ \mathbb{S}^{4k+1}\vee\mathbb{S}^{4k+1}&\mbox{if }m=6k+3\\ \mathbb{S}^{4k+2}&\mbox{if }m=6k+5\end{array}\right.$.
From this formula and Theorem 10 we get the following corollary.
@@
[…]
@@
$\mathrm{conn}(I(G\times P_{3r}))\geq 2r-1\;\;\mbox{and}\;\;\mathrm{conn}(I(G\times P_{3r+2}))\geq 2r+1.$
@@
For $n=1$, this is clear. If we show the theorem is true for $n=2$, then, by Theorem 10, it is true for any $n\geq 3$. Assume $n=2$ and let $T$ be a spanning tree of $G^{c}$. For $i=1,2$, we define $T_{i}$ as the copy of $T$ in $\Delta^{V(G)\times\{i\}}$ and we take $x$ a vertex in $V(G)$. We take
@@
Then $\pi_{1}\left(I(G\times P_{2}),(x,1)\right)\cong H_{\mathcal{T}}$ (see [rotmantop, Theorem 7.34]). We will see that $H_{\mathcal{T}}$ is trivial. Take $e$ an edge of $(G\times P_{2})^{c}$:
@@
  • If $e=\{(w,1),(w,2)\}$, we take $y_{0}y_{1}\cdots y_{r}$ the only path form $w$ to $x$ in $T_{1}$. For all $i$, we have that $\{(y_{i},1),(y_{i},2),(y_{i+1},2)\}$ is an independent set. Thus
@@
the rest of possibilities are analogous. Therefore $H_{\mathcal{T}}$ is trivial. ∎
The lower bounds given in last Theorem are tight, to see this we take $C_{5}\times P_{2}\cong C_{10}$, by Theorem 6 $I(C_{5}\times P_{2})\simeq\mathbb{S}^{2}$.
@@
For $m=1$, we have that $P_{n}\times P_{1}\cong K_{n}^{c}$ and clearly $K_{n}^{c}$ is in $\mathcal{SW}$. For $m=2$, we have that $P_{n}\times P_{2}\cong 2P_{n}$. Then, by Proposition 7, $P_{n}\times P_{2}$ is in $\mathcal{SW}$.
For $m=3$, we know the result is true for $n\leq 2$. Assume that $P_{n}\times P_{3}$ is in $\mathcal{SW}$. We take $H$ an induced subgraph of $P_{n+1}\times P_{3}$. If neither of the vertices $(1,1),(1,2),(1,3)$ are in $V(H)$, then $H$ is an induced subgraph of $P_{n+1}\times P_{3}-(1,1)-(1,2)-(1,3)\cong P_{n}\times P_{3}$ and $I(H)$ is contractible or it has the homotopy type of a wedge of spheres.
@@
Regardless of the case, we got that $I(H)$ is either contractible or $I(H)$ has the homotopy type of the double suspension of a complex $I(H^{\prime})$ where $H^{\prime}$ is an induced subgraph of $P_{n+1}\times P_{3}-(1,1)-(1,2)-(1,3)\cong P_{n}\times P_{3}$. If $(1,3)$ is in $H$, we proceed as before. If $(1,2)$ is a vertex of $H$ and $(1,1),(1,3)$ are not in $V(H)$, we proceed as the last case when we assumed $(1,2)$ was a vertex of $H-N_{H}[(2,2)]$.
Lastly, assume $m=4$. We know that the theorem is true for $n\leq 3$. Assume $P_{n}\times P_{4}$ is in $\mathcal{SW}$. We take $H$ an induced subgraph of $P_{n+1}\times P_{4}$. If neither of the vertices $(1,1),(1,2),(1,3),(1,4)$ are in $V(H)$, then $H$ is an induced subgraph of $P_{n+1}\times P_{4}-(1,1)-(1,2)-(1,3)-(1,4)\cong P_{n}\times P_{4}$ and there is nothing to prove. If the vertex $(1,1)$ is in $H$, it has at most degree $1$, then $I(H)$ is contractible or $I(H)\simeq\Sigma I(H-N_{H}[(2,2)])$. If $(1,4)$ is in $H$, then $d_{H-N_{H}[(2,2)]}((1,4))=d_{H}((1,4))\leq 1$ and $I(H-N_{H}[(2,2)])$ is contractible or
@@
with $H-N_{H}[(2,2)]-N_{H}[(2,3)]$ an induced subgraph of $P_{n+1}\times P_{4}-(1,1)-(1,2)-(1,3)-(1,4)\cong P_{n}\times P_{4}$. Assume $(1,4)$ is not in $H$. If $(2,4)$ is a vertex of $H$, then $I(H-N_{H}[(2,2)])$ is contractible. Assume $(2,4)$ is not in $H$. If $(1,2)$ is not in $H$, then $H-N_{H}[(2,2)]$ is an induced subgraph of $P_{n+1}\times P_{4}-(1,1)-(1,2)-(1,3)-(1,4)\cong P_{n}\times P_{4}$. Assume $(1,2)$ is a vertex of $H$, then $d_{H-N_{H}[(2,2)]}((1,2))=d_{H}((1,2))\leq 2$. If $d_{H}((1,2))=0$, then $I(H-N_{H}[(2,2)])$ is contractible. If $d_{H}((1,2))=1$, then
@@
Assume $(3,2)$ is not a vertex of $H$, then
@@
In either case $I(H)$ is contractible or has the homotopy type of the double suspension of a complex $I(H^{\prime})$ where $H^{\prime}$ is an induced subgraph of $P_{n+1}\times P_{4}-(1,1)-(1,2)-(1,3)-(1,4)\cong P_{n}\times P_{4}$. If $(1,4)$ is in $H$, we proceed as before. If $(1,2)$ or $(1,3)$ are vertices of $H$ and $(1,1),(1,4)$ are not in $V(H)$, we proceed as the last case when we assumed $(1,2)$ was a vertex of $H-N_{H}[(2,2)]$. ∎
Question 18.
@@
Before we conclude this section, we want to point out that while the homotopy type of the independence complex of the categorical product of paths was easy to determine, the homotopy type of $I(P_{n}\oblong P_{m})$ is only know for $1\leq m\leq 6$ (see [matsushitan4n5, matsushitan6]).